### v1

@@ Abstract
Tuberculosis (TB) remains a formidable global health challenge, driven by complex spatiotemporal transmission dynamics and influenced by factors such as population mobility and behavioral changes. We propose an Epidemic-Guided Deep Learning (EGDL) approach that fuses mechanistic epidemiological principles with advanced deep learning techniques to enhance early warning systems and intervention strategies for TB outbreaks. Our framework is built upon a networked Susceptible-Infectious-Recovered (SIR) model augmented with a saturated incidence rate and graph Laplacian diffusion, capturing both long-term transmission dynamics and region-specific population mobility patterns. Compartmental model parameters are rigorously estimated using Bayesian inference via the Markov Chain Monte Carlo (MCMC) approach. Theoretical analysis leveraging the comparison principle and Green’s formula establishes global stability properties of the disease-free and endemic equilibria. Building on these epidemiological insights, we design two forecasting architectures, EGDL-Parallel and EGDL-Series, that integrate the mechanistic outputs of the networked SIR model within deep neural networks. This integration mitigates the overfitting risks commonly encountered in data-driven methods and filters out noise inherent in surveillance data, resulting in reliable forecasts of real-world epidemic trends. Experiments conducted on TB incidence data from 47 prefectures in Japan demonstrate that our approach delivers robust and accurate predictions across multiple time horizons (short to medium-term forecasts). Additionally, incorporating uncertainty quantification through conformal prediction enhances the model’s practical utility for guiding targeted public health interventions.
@@ 1 Introduction
Tuberculosis (TB), caused by Mycobacterium tuberculosis complex, is a chronic infectious disease primarily transmitted through aerosols produced by coughing [1]. In 2023, the World Health Organization (WHO) reported 8.2 million new TB cases globally, a significant increase from 7.5 million in 2022111https://www.who.int/teams/global-tuberculosis-programme/tb-reports. This makes TB the leading cause of death from a single infectious disease, surpassing COVID-19. To address the global health burden associated with a high TB incidence rate, the WHO introduced the “End TB strategy” in 2015, aiming to reduce TB incidence cases below 0.01% by 2035 and 0.001% by 2050 [2]. However, the current global rate of decline is not on track to achieve these goals. In the era of rapid globalization, the existence of a TB-free region is impractical unless a global mitigation approach is successfully implemented. While low and medium-income countries are traditionally considered high TB-burden regions, TB remains a persistent challenge even in high-income countries due to increased population mobility [3]. Among the high-income nations, Japan continues to be classified as a TB middle-burden country, with an incidence rate of 10-49 cases per 100,000 population [^2] https://iris.who.int/handle/10665/341980. High TB incidences and related mortality in Japan are primarily attributed to an aging population [4], as well as the rising number of immigrants from high TB-burden countries, accounting for a growing proportion of cases among younger populations [5]. To mitigate the health and socioeconomic consequences of TB, various attempts have been made to develop early warning systems that predict the future TB dynamics [6, 7]. An early warning system provides quick hindsight, allowing for timely intervention and better disease control in its nascent stages. Epidemic modeling and forecasting are central to these efforts, as they enable more accurate predictions of disease trends, guide public health responses, and help allocate resources more effectively to reduce the TB transmission rate.
@@
Compartmental epidemiological models are widely used to characterize infectious disease dynamics by dividing the population into distinct and mutually exclusive subsets based on their infection status [8, 9]. The compartmental SIR model categorizes individuals as Susceptible ($S$), those who are vulnerable to infection; Infectious ($I$), those who are currently infected and actively transmitting the disease; and Recovered ($R$), those who have gained immunity after recovering from the disease [10]. Numerous extensions of the classical SIR model, including additional compartments and refined incidence functions, have been developed to model TB transmission dynamics in Indonesia [11] and Kazakhastan [12]. Although compartmental epidemiological models yield fundamental insights into infectious disease outbreaks, they are more suited for understanding the disease dynamics, rather than real-time forecasting of the disease incidences [13]. To overcome the problem of limited predictability of the compartmental approaches, data-centric forecasting methods, including statistical and deep learning models, have emerged as powerful tools. For example, [14] demonstrated the effectiveness of deep learning frameworks in forecasting TB infections in Brazil, while [7] implemented a data-driven hybrid framework to predict TB incidence in China. These approaches leverage diverse data sources to uncover hidden patterns in epidemic data, showcasing their flexibility. However, this flexibility can often be a double-edged sword, making them prone to overfitting, particularly in low-data regimes, a common challenge in epidemic forecasting [15]. Moreover, these models often generate unrealistic predictions for non-stationary systems as they solely rely on lagged values of disease incidences. The primary disadvantage of data-driven methods for epidemic forecasting is that they are not guided by the mechanisms governing disease transmission. By focusing solely on the dynamics of surveillance data without integrating epidemiological principles, these approaches often generate unrealistic forecasts and struggle to differentiate genuine trends from noise introduced during data collection [16]. To bridge this gap, a recent line of research has explored integrating epidemic dynamics from compartmental models into data-centric forecasting frameworks [17, 18]. For instance, Epidemic-Informed Neural Networks (EINNs) embed epidemiological knowledge from ordinary differential equation (ODE) based compartmental models into Recurrent Neural Networks (RNN) via gradient matching, enhancing model flexibility and resilience against noisy data [19]. Similarly, [16] employed Physics Informed Neural Networks (PINNs) to incorporate epidemiological principles from a modified SIR framework with observed infectious data, improving forecast accuracy through transfer learning.
@@
Despite these advancements, most studies in this domain rely on traditional SIR-based compartmental models, which do not account for spatial dynamics, such as population mobility. As a result, these models often fail to capture the spatiotemporal trajectory of an epidemic, limiting their ability to predict real-world transmission patterns accurately. Recent epidemiological studies have incorporated network-based structures into SIR models to address this limitation in compartmental models, enhancing their ability to capture spatial disease transmission [20, 21]. By leveraging network structures, the SIR model can estimate epidemic dynamics at the population level rather than focusing solely on individual-level infection status, leading to a more comprehensive understanding of disease spread. To represent the spatial disease dynamics, previous studies have often relied on Laplacian operators to model population mobility, assuming isotropic diffusion and uniform movement probabilities [22, 23, 24]. However, these assumptions are frequently violated in real-world scenarios, where human mobility is influenced by local environmental factors such as lockdowns, travel restrictions, and other region-specific circumstances. To address these challenges, recent studies have extended the global stability theory of the SIR model by incorporating graph Laplacian diffusion, leading to more efficient frameworks for modeling spatial epidemic dynamics [25, 26]. In existing networked epidemic models, bilinear incidence rate ($\beta SI$) and standard bilinear incidence rate ($\beta SI/N$), with transmission rate $\beta$ and population size $N$, are widely used to analyze the global dynamics of disease spread [27, 28]. However, these incidence functions fail to account for behavioral changes and crowding effects, often leading to overestimated infection rates in highly connected populations. To overcome these limitations, we adopt a saturated incidence function [29] and formulate a networked SIR model by incorporating graph Laplacian diffusion with the saturated incidence rate. This integration better captures long-term transmission dynamics and provides an adaptive framework for modeling dynamic changes in population behavior. Building on this epidemiological foundation, we further introduce the Epidemic-Guided Deep Learning (EGDL) approach, which combines spatial epidemiological insights with data-driven forecasting techniques. EGDL generates accurate epidemic forecasts by considering both spatial dynamics and temporal interactions. The proposed EGDL architectures integrate the long-term infection dynamics of the networked SIR model as auxiliary information within deep learning frameworks. This integration combines the flexibility of deep learning techniques and the epidemiological principles captured by the modified version of the networked SIR model, improving spatiotemporal disease incidence forecasting. Furthermore, using epidemic dynamics as auxiliary information, instead of solely relying on gradient-based learning, restricts the overfitting risks in EGDL frameworks, allowing the model to learn robust representations. This effectively filters out noise from historical incidence data and aligns spatiotemporal disease forecasts with real-world epidemic trends. The key contributions of the study can be summarized as follows:
@@
  • We introduce a modified networked SIR model by incorporating a saturated incidence rate and graph Laplacian diffusion to understand epidemic data’s spatiotemporal dynamics. We validate the key epidemiological properties, such as positivity and boundedness. The global stability of disease-free and endemic equilibria is established using Green’s formula and the comparison principle.
@@
  • To enhance spatiotemporal disease forecasting, we propose the EGDL-Parallel and EGDL-Series architectures, combining the networked SIR model with the historical surveillance data-driven deep learning frameworks (e.g., Transformers, NBeats).
@@
  • The forecasting performance of the EGDL architectures is evaluated using a rolling window approach across four test horizons and compared against data-centric models, with robustness validated through non-parametric statistical tests. The uncertainty of the EGDL frameworks is quantified using the conformal prediction approach.
2 Motivating Example: Prefecture-specific active TB study
@@
Figure 1: (A) Geographic distribution of Japan’s 47 prefectures, shown for illustrative purposes only, without implying any political assertions on Japan’s territorial boundaries. (B) Monthly active tuberculosis (TB) cases recorded across Japan’s 47 prefectures over 216 months, spanning from January 1998 to December 2015.
TB remains a leading cause of infectious disease-related mortality, presenting a significant global public health challenge. In Japan, classified as a TB medium-burden country, the TB surveillance center reported 11,519 new cases in 2022333https://jata-ekigaku.jp/english/tb-in-japan. To control the rapid spread of TB, Japan has established a robust nationwide surveillance system that collects detailed TB data at the prefectural level, ensuring comprehensive coverage. Japan’s administrative structure, comprising 47 prefectures further divided into cities, wards, and blocks, facilitates precise data collection and monitoring. This hierarchical organization enables tracking TB trends both geographically and temporally, enhancing the effectiveness of public health interventions. In this study, we analyze the monthly counts of newly registered active TB cases of all forms across Japan’s 47 prefectures, as illustrated in Fig. 1(A). The dataset, spanning 216 months from January 1998 to December 2015, is publicly available at [30, 31]. Fig. 1(B) visually represents the monthly active TB cases for all 47 prefectures. In this figure, each prefecture is represented by a distinct color, enabling clear differentiation and facilitating visual comparisons of TB trends across regions. These TB datasets represent a spatiotemporal forecasting problem for infectious disease modelers. To capture the spatiotemporal dynamics of TB outbreaks, our study aims to build an integrated solution based on epidemiological models and deep learning approaches.
@@ 3 Preliminaries
Epidemic modeling and forecasting approaches can be broadly categorized as mechanistic (or compartmental) frameworks and phenomenological models. Mechanistic models use causal frameworks to describe disease states and understand the surveillance data-generating process. In contrast, phenomenological models focus on directly modeling surveillance data (past lagged observations of the epidemic time series) without explicitly incorporating the underlying epidemiological mechanisms. The following subsections provide a brief overview of the mechanistic frameworks (Section 3.1) and phenomenological models (Section 3.2) to be used as building blocks for the proposed EGDL frameworks.
@@ 3.1 Epidemic Models: A Review
where $\alpha$ represents the saturation factor. This model assumes homogeneous mixing among individuals, meaning that every susceptible individual is equally likely to interact with any other individual in the population at any given time. This assumption simplifies the representation of contact patterns, treating all interactions as uniformly distributed across the population. In real-world population dynamics, contact patterns are often heterogeneous rather than homogeneous. Building upon this basic framework, heterogeneity can be introduced by further subdividing the compartments to account for diverse population characteristics [33]. Typically, the number of contacts each individual has is significantly smaller than the total population size. In such cases, the homogeneous-mixing assumption becomes inefficient. The simple SIR model describes the temporal dynamics of an infectious disease in a population; however, it fails to capture the spatial dynamics of disease incidence. Epidemic models incorporating network structures address this limitation by assigning each individual a finite set of permanent contacts within a localized region, providing a more accurate depiction of infection dynamics.

### v2

@@ Abstract
Tuberculosis (TB) remains a formidable global health challenge, driven by complex spatiotemporal transmission dynamics and influenced by factors such as population mobility and behavioral changes. We propose an Epidemic-Guided Deep Learning (EGDL) approach that fuses mechanistic epidemiological principles with advanced deep learning techniques to enhance early warning systems and intervention strategies for TB outbreaks. Our framework is built upon a modified networked Susceptible-Infectious-Recovered (MN-SIR) model augmented with a saturated incidence rate and graph Laplacian diffusion, capturing both long-term transmission dynamics and region-specific population mobility patterns. Compartmental model parameters are rigorously estimated using Bayesian inference via the Markov Chain Monte Carlo approach. Theoretical analysis leveraging the comparison principle and Green’s formula establishes global stability properties of the disease-free and endemic equilibria. Building on these epidemiological insights, we design two forecasting architectures, EGDL-Parallel and EGDL-Series, that integrate the mechanistic outputs of the MN-SIR model within deep neural networks. This integration mitigates the overfitting risks commonly encountered in data-driven methods and filters out noise inherent in surveillance data, resulting in reliable forecasts of real-world epidemic trends. Experiments conducted on TB incidence data from 47 prefectures in Japan and 31 provinces in mainland China demonstrate that our approach delivers robust and accurate predictions across multiple time horizons (short to medium-term forecasts), supporting its generalizability across regions with different population dynamics. Additionally, incorporating uncertainty quantification through conformal prediction and explainability via temporal gradient-based class activation maps enhances the model’s practical utility for guiding targeted public health interventions.
@@ 1 Introduction
Tuberculosis (TB), caused by Mycobacterium tuberculosis complex, is a chronic infectious disease primarily transmitted through aerosols produced by coughing [1]. In 2023, the World Health Organization (WHO) reported 8.2 million new TB cases globally, a significant increase from 7.5 million in 2022111https://www.who.int/teams/global-tuberculosis-programme/tb-reports. This makes TB the leading cause of death from a single infectious disease, surpassing COVID-19. To support the Sustainable Development Goal 3: “Good Health and Well-being” and to address the global health burden of high TB incidence rate, the WHO introduced the “End TB strategy” in 2015. This initiative aims to reduce TB incidence cases below 0.01% by 2035 and 0.001% by 2050 [2]. However, the current global rate of decline is not on track to achieve these goals. In the era of rapid globalization, the existence of a TB-free region is impractical unless a global mitigation approach is successfully implemented. While low and medium-income countries are traditionally considered high TB-burden regions, TB remains a persistent challenge even in high-income countries due to increased population mobility [3]. Among the high-income nations, Japan continues to be classified as a TB middle-burden country, with an incidence rate of 10-49 cases per 100,000 population [^2] https://iris.who.int/handle/10665/341980. High TB incidences and related mortality in Japan are primarily attributed to an aging population [4], as well as the rising number of immigrants from high TB-burden countries, accounting for a growing proportion of cases among younger populations [5]. Alongside Japan, mainland China remains one of the countries with a high-burden of TB incidences owing to its large population, latent infections, and disparities in healthcare access across regions [6]. To mitigate the health and socio-economic consequences of TB, various attempts have been made to develop early warning systems that predict the future TB dynamics [7, 8]. An early warning system provides quick hindsight, allowing for timely intervention and better disease control in its nascent stages. Epidemic modeling and forecasting are central to these efforts, as they enable more accurate predictions of disease trends, guide public health responses, and help allocate resources more effectively to reduce the TB transmission rate.
@@
Compartmental epidemiological models are widely used to characterize infectious disease dynamics by dividing the population into distinct and mutually exclusive subsets based on their infection status [9, 10]. The compartmental SIR model categorizes individuals as Susceptible ($S$), those who are vulnerable to infection; Infectious ($I$), those who are currently infected and actively transmitting the disease; and Recovered ($R$), those who have gained immunity after recovering from the disease [11]. Numerous extensions of the classical SIR model, including additional compartments and refined incidence functions, have been developed to model TB transmission dynamics in Indonesia [12] and Kazakhstan [13]. Although compartmental epidemiological models yield fundamental insights into infectious disease outbreaks, they are more suited for understanding the disease dynamics, rather than real-time forecasting of the disease incidences [14, 15]. To overcome the problem of limited predictability of the compartmental approaches, data-centric forecasting methods, including statistical and deep learning models, have emerged as powerful tools. For example, [16] demonstrated the effectiveness of deep learning frameworks in forecasting TB infections in Brazil, while [8] implemented a data-driven hybrid framework to predict TB incidence in China. These approaches leverage diverse data sources to uncover hidden patterns in epidemic data, showcasing their flexibility. However, this flexibility can often be a double-edged sword, making them prone to overfitting, particularly in low-data regimes, a common challenge in epidemic forecasting [17]. Additionally, these models often generate unrealistic predictions for non-stationary systems as they solely rely on lagged values of disease incidences. The primary disadvantage of data-driven methods for epidemic forecasting is that they are not guided by the mechanisms governing disease transmission. By focusing solely on the dynamics of surveillance data without integrating epidemiological principles, these approaches often generate unrealistic forecasts and struggle to differentiate genuine trends from noise introduced during data collection [18]. To bridge this gap, a recent line of research has explored combining epidemic dynamics from compartmental models into data-centric forecasting frameworks [19, 20]. For instance, Epidemic-Informed Neural Networks (EINNs) embed epidemiological knowledge from ordinary differential equation (ODE) based compartmental models into Recurrent Neural Networks (RNN) via gradient matching, enhancing model flexibility and resilience against noisy data [21]. Similarly, [18] employed Physics Informed Neural Networks (PINNs) to incorporate epidemiological principles from a modified SIR framework with observed infectious data, improving forecast accuracy through transfer learning.
@@
Despite these advancements, most studies in this domain rely on traditional SIR-based compartmental models, which do not account for spatial dynamics, such as population mobility. As a result, these models often fail to capture the spatiotemporal trajectory of an epidemic, limiting their ability to predict real-world transmission patterns accurately. Recent epidemiological studies have incorporated network-based structures into SIR models to address this limitation in compartmental models, enhancing their ability to capture spatial disease transmission [22, 23]. By leveraging network structures, the SIR model can estimate epidemic dynamics at the population level rather than focusing solely on individual-level infection status, leading to a more comprehensive understanding of disease spread. To represent the spatial disease dynamics, previous studies have often relied on Laplacian operators to model population mobility, assuming isotropic diffusion and uniform movement probabilities [24, 25, 26]. However, these assumptions are frequently violated in real-world scenarios, where human mobility is influenced by local environmental factors such as lockdowns, travel restrictions, and other region-specific circumstances. To address these challenges, recent studies have extended the global stability theory of the SIR model by incorporating graph Laplacian diffusion, leading to more efficient frameworks for modeling spatial epidemic dynamics [27, 28]. In existing networked epidemic models, bilinear incidence rate ($\beta SI$) and standard bilinear incidence rate ($\beta SI/N$), with transmission rate $\beta$ and population size $N$, are widely used to analyze the global dynamics of disease spread [29, 30]. However, these incidence functions fail to account for behavioral changes and crowding effects, often leading to overestimated infection rates in highly connected populations. To overcome these limitations, we adopt a saturated incidence function [31] and formulate a networked SIR model by incorporating graph Laplacian diffusion, we name it modified networked SIR (MN-SIR) model. This integration better captures long-term transmission dynamics and provides an adaptive framework for modeling dynamic changes in population behavior. Building on this epidemiological foundation, we further introduce the Epidemic-Guided Deep Learning (EGDL) approach, which combines spatial epidemiological insights with data-driven forecasting techniques. EGDL generates accurate epidemic forecasts by considering both spatial dynamics and temporal interactions. The proposed EGDL architectures integrate the long-term infection dynamics of the MN-SIR model as auxiliary information within deep learning frameworks. This integration combines the flexibility of deep learning techniques and the epidemiological principles captured by the MN-SIR model, improving spatiotemporal disease incidence forecasting. Furthermore, using epidemic dynamics as auxiliary information, instead of solely relying on gradient-based learning, restricts the overfitting risks in EGDL frameworks, allowing the model to learn robust representations. This effectively filters out noise from historical incidence data and aligns spatiotemporal disease forecasts with real-world epidemic trends. The key contributions of the study can be summarized as follows:
@@
  • We introduce a MN-SIR model by incorporating a saturated incidence rate and graph Laplacian diffusion to understand epidemic data’s spatiotemporal dynamics. We validate the key epidemiological properties, such as positivity and boundedness. The global stability of disease-free and endemic equilibria is established using Green’s formula and the comparison principle.
@@
  • To enhance spatiotemporal disease forecasting, we propose the EGDL-Parallel and EGDL-Series architectures, combining the MN-SIR model with the historical surveillance data-driven deep learning frameworks (e.g., Transformers, NBeats).
@@
  • The forecasting performance of the EGDL architectures is evaluated using real-world TB incidence datasets from 47 prefectures of Japan and 31 provinces of mainland China via a rolling window approach across four test horizons. We compared the performance the EGDL frameworks against data-centric models and validated the robustness through a non-parametric statistical test. The uncertainty of the EGDL frameworks is quantified using the conformal prediction approach, while the explainability of the model output is obtained via temporal gradient-based class activation maps.
2 Motivating Examples
@@
Figure 1: (A) Geographic distribution of Japan’s 47 prefectures, shown for illustrative purposes only, without implying any political assertions on Japan’s territorial boundaries. (B) Monthly active tuberculosis (TB) cases were recorded in each of Japan’s 47 prefectures from January 1998 to December 2015.
TB remains one of the leading causes of infectious disease-related mortality, posing a significant global public health challenge. This study focuses on TB incidence across two major Asian countries, Japan and mainland China, which represent contrasting TB burden levels and demographic profiles. Japan, classified as a TB medium-burden country, reported 11,519 new TB cases in 2022333https://jata-ekigaku.jp/english/tb-in-japan. To control the rapid spread of TB, Japan has established a robust nationwide surveillance system that collects granular data at the prefectural level, facilitated by the country’s administrative structure of 47 prefectures. This hierarchical setup enables precise geographic and temporal monitoring of TB trends, enabling timely public health responses. In this study, we analyze the monthly counts of newly registered active TB cases (of all forms) across Japan’s 47 prefectures, as illustrated in Fig. 1(A). The dataset spans from January 1998 to December 2015 and is publicly available at [32, 33]. Fig. 1(B) displays monthly TB cases for each prefecture, with distinct colors to facilitate regional comparison. In addition to Japan, we analyze TB incidence in mainland China, a high-burden country that ranked third globally in estimated TB cases in 2022444TB Disease Burden (WHO).. Despite steady progress in TB prevention, the disease remains a major public health concern due to China’s large population and the widespread prevalence of latent infections. Consequently, TB is classified as a Class B respiratory infectious disease in China, necessitating ongoing monitoring [6]. Additionally, the COVID-19 pandemic severely disrupted healthcare systems, affecting the continuity and reliability of TB surveillance. To ensure data consistency, we restrict our analysis to the pre-pandemic period, focusing on monthly notified TB cases from 31 provinces between January 2014 and December 2018, which were collected from the Chinese Center for Disease Control and Prevention [^5] https://en.chinacdc.cn/ [6]. Fig. 2 displays the TB incidence trends across these provinces during the study period, with each province uniquely color-coded to highlight regional variations. Together, the TB datasets from Japan and China offer complementary insights into disease dynamics across East Asia, ranging from a medium-burden, aging society to a high-burden, populous nation facing latent infections. This regional heterogeneity presents a challenging spatiotemporal forecasting problem. Our study aims to solve this problem by developing an integrated solution, based on epidemiological models and deep learning approaches, that can accurately predict future TB transmission patterns.
Figure 2: (A) Geographic distribution of 31 provinces of mainland China, shown for illustrative purposes only, without implying any political assertions on China’s territorial boundaries. (B) Log-transformed values of monthly active tuberculosis (TB) cases recorded in each of China’s 31 provinces from January 2014 to December 2018 (log transformation is applied only for visual clarity).
@@ 3 Preliminaries
Epidemic modeling and forecasting approaches can be broadly categorized as mechanistic (or compartmental) frameworks and phenomenological models. Mechanistic models use causal frameworks to describe disease states and understand the surveillance data-generating process. In contrast, phenomenological models focus on directly modeling surveillance data (past lagged observations of the epidemic time series) without explicitly incorporating the underlying epidemiological mechanisms. The following subsections provide a brief overview of the mechanistic frameworks (Section 3.1), spatiotemporal forecasting task (Section 3.2), and phenomenological deep learning models (Section 3.3) to be used as building blocks for the proposed EGDL frameworks.
@@ 3.1 Epidemic Models: A Review
where $\alpha$ represents the saturation factor. We assume that $\Lambda$ represents the constant recruitment or birth rate into the susceptible population at each node. The parameter $\mu$ is the natural death rate, included in all compartments to model uniform mortality. The term $(\gamma+\mu)I$ accounts for the removal of infected individuals due to recovery ($\gamma$) or death ($\mu$), while $\gamma I$ contributes to the recovered class. This model also assumes homogeneous mixing among individuals, meaning that every susceptible individual is equally likely to interact with any other individual in the population at any given time. This assumption simplifies the representation of contact patterns, treating all interactions as uniformly distributed across the population. In real-world population dynamics, contact patterns are often heterogeneous rather than homogeneous. Building upon this basic framework, heterogeneity can be introduced by further subdividing the compartments to account for diverse population characteristics [35]. Typically, the number of contacts each individual has is significantly smaller than the total population size. In such cases, the homogeneous-mixing assumption becomes inefficient. The simple SIR model describes the temporal dynamics of an infectious disease in a population; however, it fails to capture the spatial dynamics of disease incidence. Epidemic models incorporating network structures address this limitation by assigning each individual a finite set of permanent contacts within a localized region, providing a more accurate depiction of infection dynamics.
3.2 Spatiotemporal Forecasting: Problem Definition
Time series data consists of a sequence of observations recorded at uniformly spaced time intervals [36]. Based on the number of variables measured at each time step, temporal datasets are broadly categorized into univariate and multivariate time series. A univariate time series contains a sequence of scalar observations collected over time, whereas a multivariate time series comprises sequences of vector-valued observations, where multiple variables are recorded simultaneously. Let $V=\{1,2,\dots,n\}$ denote a set of spatial locations (e.g., prefectures or provinces), and let $\mathcal{T}=\{1,2,\dots,T\}$ denote discrete and equally spaced time points. A univariate time series at location $x\in V$ is a sequence $\{Y(x,t)\}_{t\in\mathcal{T}}$, where $Y(x,t)\in\mathbb{R}_{\geq 0}$ represents the observed number of TB cases at location $x$ and time $t$. A multivariate time series refers to a sequence of vectors $\{\mathbf{y}_{t}\}_{t=1}^{T}$, where $\mathbf{y}_{t}\in\mathbb{R}^{d}$ encodes $d$-dimensional observations at time $t$. In our setup, this corresponds to stacking TB incidence values from all locations at time $t$, i.e., $\mathbf{y}_{t}=[Y(1,t),Y(2,t),\dots,Y(n,t)]^{\top}\in\mathbb{R}^{n}$, where $d=n$ is the number of spatial units. With these multiple variables observed across different spatial locations over time, the multivariate time series naturally extends to a spatiotemporal dataset. The collection $\{Y(x,t):x\in V,t\in\mathcal{T}\}$ forms a spatiotemporal time series, which captures both temporal dynamics (evolution over time) and spatial interactions (across locations) [37]. One of the primary tasks in time series analysis is forecasting, which involves predicting future values based on historical observations. Depending on the nature of the temporal data, forecasting can be categorized into univariate and multivariate forecasting. Univariate forecasting focuses on predicting future values of a single time series using only its past observations. In contrast, multivariate forecasting predicts target variables by leveraging historical data from multiple related time series, capturing the inter-dependencies among correlated variables and often improving forecast accuracy.
Spatiotemporal forecasting further extends the multivariate forecasting task by jointly modeling multiple interrelated time series distributed across both space and time, thereby capturing complex spatial and temporal patterns. The multi-step ahead forecasting task is defined as follows: given historical data up to time $T$, i.e., $\{Y(x,t)\}_{x\in V,t\leq T}$, the objective is to predict future values $\{\widehat{Y}(x,T+v)\}_{v=1}^{q}$ for each location $x\in V$, where $q\in\{1,2,\ldots\}$ is the forecasting horizon. This is a multi-output forecasting problem, involving the simultaneous prediction of future values across multiple spatial units and time steps [38]. Let $f_{M}\in\mathcal{F}$ be a forecasting model, where $\mathcal{F}$ denotes the class of data-driven global forecasting architectures (for e.g. Transformers, TCN, NBeats, and NHits). The model $f_{M}$ takes as input a sequence of past observations over a rolling window of size $t_{w}$ and generates forecasts for the next $q$ time steps. Thus using the historical data $\{Y(x,t)\}_{x\in V,t\leq T}$ for training $f_{M}$ the future dynamics can be predicted as follows:
where $x\in V,\;\tilde{t}=t_{w},t_{w}+1,\ldots,T,\;v=1,2,\ldots,q$, and $t_{w}$ denotes the number of historical inputs used for predicting the subsequent observations. Note that the above formulation is consistent with the state-of-the-art multivariate forecasting models and also generalizable to spatiotemporal variants considered in this study.